f^(7) (x) of y = x^-2


1
Expert's answer
2021-10-31T18:30:05-0400

Let "P(n)" be the proposition that "f^{(n)}=(-1)^{n}(n+1)!x^{-(n+2)}, n\geq1"

Basis Step

"P(1)" is true, because


"f'(x)=(x^{-2})'=-2x^{-3}=(-1)^1(1+1)!x^{-(1+2)}"

Inductive Step

Assume that "P(k)" holds for an arbitrary integer "k\geq1." That is, we assume that


"f^{(k)}=(-1)^{k}(k+1)!x^{-(k+2)}, k\geq1"

Under this assumption, it must be shown that "P(k + 1)" is true, namely, that


"f^{(k+1)}=(-1)^{k+1+1}(k+1)!x^{-(k+1+2)},"

is also true.


"f^{(k+1)}=(f^{(k)})'=((-1)^{k}(k+1)!x^{-(k+2)})'"

"=(-1)^{k}(k+1)!(-(k+2))x^{-(k+2)-1}"

"=(-1)^{k+1}(k+2)!x^{-(k+1+2)}"

"=(-1)^{k+1}(k+1+1)!x^{-(k+1+2)}"

This last equation shows that "P(k + 1)" is true under the assumption that "P(k)" is true. This completes the inductive step.

We have completed the basis step and the inductive step, so by mathematical induction we

know that "P(n)" is true for all integers "n\geq1." That is, we have proven that 


"f^{(n)}=(-1)^{n}(n+1)!x^{-(n+2)}, n\geq1"

Then


"f^{(7)}=(-1)^{7}(7+1)!x^{-(7+2)}=-8!x^{-9}"

"=-40320x^{-9}"